# Scalable analysis of stop-and-go waves

### v1

@@ 4 Methodology
$N$ the number of vehicle trajectories - -
@@
$\bm{\mathrm{\tau}}_{i}$ vector representing the trajectory of vehicle $i$
$\bm{\mathrm{\tau}}_{i}^{c}$ vector representing the stationary component trajectory of vehicle $i$
$\bm{\mathrm{\tau}}_{i}^{osc}$ vector representing the oscillation component trajectory of vehicle $i$
$\epsilon$ local search boundary for local minimum and maximum second Yes
@@
$d_{i}$ wave fronts set for vehicle $i$
$a_{i}$ wave tails set for vehicle $i$
$K_{i}$ the number of wave fronts and tails for vehicle $i$
@@
$N_{c}$ the number of connected components for graph $\mathcal{G}$
$\mathcal{C}^{d}$ the components for the wave fronts graph $\mathcal{G}_{d}$
$\mathcal{C}^{a}$ the components for the wave tails graph $\mathcal{G}_{a}$
@@ 4.1 Preliminaries
Long-range raw vehicle trajectory data as from [9] is imperfect due to the challenges of tracking and re-identification in dense traffic conditions [97; 98]. In essence, it is difficult to track vehicles perfectly across hundreds of cameras and through occlusions [99], so trajectory datasets over large spatial ranges often contain fragmented trajectories. (These issues do not significantly impact local traffic measurements, such as traffic speeds, which can still be obtained with extremely high fidelity.) We prepare raw data for the methods described next by utilizing the method from [100], which we describe in more detail in A. The resulting ’virtual trajectories’ have no passing or lane change maneuvers. This allows efficient search between adjacent trajectories, but the methods described next could also be applied to raw trajectory data with lane-change maneuvers with only slight modification.
@@ 4.2 Identifying the stop-and-go waves: the trajectory perspective
From the perspective of a vehicle trajectory, stop-and-go waves can be intuitively understood as a sequence of deceleration and acceleration events [14; 18]. This pattern is discernible from the speed time-series profile of each individual vehicle (see Figure 3). By analyzing these profiles, we can observe the fluctuations in speed that characterize stop-and-go waves, providing a clear visualization of how vehicles repeatedly slow down and speed up within the traffic flow.
@@
Let $\mathcal{T}$ be the set of all trajectories. (Note that the preprocessing in A ensures that trajectories are ordered by increasing time, and no trajectories enter or exit the considered lane). Let $\bm{\mathrm{\tau}}_{i}\in\mathcal{T}$ represent the trajectory of vehicle $i$ as a vector of time-space points. Throughout this section, we will use $\tau_{i}(t)$ to represent a continuous, functional representation of the trajectory of vehicle $i$ from which the points in $\bm{\mathrm{\tau}}_{i}$ are sampled. To detect the critical events from the trajectories, each trajectory is decomposed into two components as follows:
@@
By definition, a wave front or tail occurs when trajectory velocity $\dot{\tau_{i}(t)}=v_{c}$, or when $\dot{\tau_{i}^{\text{osc}}}(t)=0$. Thus, wave front and tail points correspond to a local maxima or minima of $\tau_{i}^{osc}(t)$, respectively.
  • Wave front points. The local maximum point $t^{\text{front}}=t_{\text{local}}^{\text{max}}$, representing the wave front, if there exists an interval $(t_{\text{local}}^{\text{max}}-\epsilon,t_{\text{local}}^{\text{max}}+\epsilon)$ such that for all $t$ in this interval (except $t_{\text{local}}^{\text{max}}$), the following condition holds:
    $\tau_{i}^{osc}(t_{\text{local}}^{\text{max}})\geq\tau_{i}^{osc}(t),$ (4)
    where $\epsilon$ denotes the local search boundary for local maximum.
  • Wave tail points. The local minimum point $t^{\text{tail}}=t_{\text{local}}^{\text{min}}$, representing the wave tail, if there exists an interval $(t_{\text{local}}^{\text{min}}-\epsilon,t_{\text{local}}^{\text{min}}+\epsilon)$ such that for all $t$ in this interval (except $t_{\text{local}}^{\text{min}}$), the following condition holds:
    $\tau_{i}^{osc}(t_{\text{local}}^{\text{min}})\leq\tau_{i}^{osc}(t),$ (5)
where $\epsilon$ denotes the local search boundary for local minimum. Using $\bm{\mathrm{\tau}}_{i}^{osc}$ as a proxy for $\tau_{i}^{osc}(t)$, we find all local maxima and minima in $\bm{\mathrm{\tau}}_{i}^{osc}$, yielding a set $d_{i}$ of wave front and a set $a_{i}$ of wave tail points for trajectory $i$. We keep only those wave fronts that are followed by a wave tail and those wave tails that are preceded by a wave front, excluding any wave fronts or tails that occur at the boundaries of the trajectories, as they do not contribute to a complete stop-and-go cycle. This ensures that there are the same number $K_{i}$ of wave front points and wave tail points for trajectory $i$. This process is repeated for all trajectories in $\mathcal{T}$, yielding an overall set of wave front and tail points $\mathcal{D}=\{d_{0},d_{1},...d_{i}\}$ and $\mathcal{A}=\{a_{0},a_{1},...a_{i}\}$.
@@ 4.3.2 Edges
Next, we describe the process for adding edges to the stop-and-go wave graph such that when finished, all connected wave fronts and wave tails correspond to a single distinct traffic wave. Two types of edges are considered to connect nodes in the stop-and-go-wave graph. i). Inner edges connect wave fronts and tails for the same trajectory, based on the intuition that each wave front and tail are added to the graph as a pair and by definition correspond to the same stop and go wave. ii.) Cross edges connect either wave fronts or wave tails for nearby trajectories, based on the idea two decelerations or accelerations experienced by two trajectories in close spatio-temporal proximity are caused by the same stop-and-go wave. This framework facilitates the analysis of both individual trajectory patterns and the interactions between different trajectories [14; 13].
@@
    $\mathcal{E}_{\text{inner}}=\bigcup_{i=1}^{N}\bigcup_{k_{i}=1}^{K_{i}}\{\{d_{i,k_{i}},a_{i,k_{i}}\}\}$ (7)
@@
    $\mathcal{E}_{\text{cross}}=\mathcal{E}_{\text{cross}}^{\mathcal{D}}\cup\mathcal{E}_{\text{cross}}^{\mathcal{A}}$ (8)
    where the connection for wave front is illustrated by the red dashed lines connecting adjacent red dots in 4b, green for wave tails. Succinctly, for each wave front point, we search in a narrow spatio-temporal box in time and in space around that point for other wave front points identified on the next upstream trajectory. If any are found, we add an edge to $\mathcal{E}_{\text{cross}}^{\mathcal{D}}$ corresponding to the closest of these points (in time). The same is then done for each wave tail point. Mathematically, we define these sets of edges connecting wave fronts ($\mathcal{E}_{\text{cross}}^{\mathcal{D}}$) and tails ($\mathcal{E}_{\text{cross}}^{\mathcal{A}}$) as:
@@ 5 Data
I-24 MOTION is a traffic instrument for the freeway traffic observation [104; 9; 98], which is designed for the continuous freeway traffic data collection and analysis. It is expected to generate 200 million vehicle miles traveled (VMT) in trajectory data annually [98], which poses significant challenges for data analysis. In this paper, the I-24 MOTION INCEPTION data released by [9] is used for the stop-and-go wave analysis with the proposed method in Section 4. Table 4 summarizes the data used in this paper, including the corresponding week day of the date, vehicle miles traveled (VMT) and the mean speed (MS).
@@
The hyper-parameter in this paper is the critical speed $v_{c}$ and is examined at various values: $\{1,5,10,15,20,25,30,35,40\}$ mph. The search region $\mathcal{S}$ as illustrated in Figure 4, is defined as a rectangular area centered around a given spatio-temporal feature, extending from -0.02 miles to 0.05 miles spatially and from -5 seconds to 15 seconds temporally. In this context, the negative direction in space refers to the opposite direction of traffic flow, while negative values in time indicate past moments relative to the spatio-temporal feature.
6 Results
@@ 6.2 Wave fronts and tails identification
For each $v_{c}$ and each lane, a set of wave fronts and tails is generated. Figures 7(a) and 7(b) illustrate the average number of wave fronts and tails identified per day, respectively. As illustrated in the figures, the number of wave fronts and tails increases as the critical speed $v_{c}$ rises. The number of wave fronts and tails tends to be higher in lane 4, followed by lane 3, lane 2, and lane 1, for speeds ranging from 10 to 40 mph. However, as shown in Figures 7(c) and 7(d), the traveling distance of the fronts and tails follows the inverse order, indicating that wave propagation is more prolonged in the inner lanes compared to the outer lanes, which may be explained by the boundary effects of on-ramps and off-ramps.
@@
Figures 7(c) and 7(d) demonstrate that the travel distances of wave fronts and tails exhibit nonlinear variations with the critical speed $v_{c}$ across four lanes. The travel distance reaches its maximum around a critical speed of 15-20 mph for all lanes, indicating that this speed range is most critical for wave propagation. Lane 1 consistently has the longest wave front travel distance, while lane 4 shows the shortest across all speed levels. These findings suggest that there is a critical speed range for maximizing wave propagation distance, with notable differences in wave dynamics between lanes, particularly at lower and higher speeds. This emphasizes the sensitivity of wave behavior to both lane-specific factors and the chosen critical speed.
@@ 6.3 Wave components identification
7 Discussion
@@ 7.2 Differences in traveling speeds between wave fronts and tails
Based on the definitions in Equations (14) and (15), each sample (i.e., wave front or tail) can be represented in a scatter plot illustrating the relationship between wave travel speed and travel distance, as shown in Figure 9. Note that the speed shown here are absolute values; in reality, these speeds represent the traveling waves in the opposite direction of traffic flow. Figure 9 illustrates the centrality of wave speed for both the fronts and tails as the wave travel distance increases at the critical speed 15 mph. The wave travel speed for distances below 0.5 miles is highly random across all lanes. Additional plots for different critical speeds can be found in C.
@@ 7.3 Wave bifurcation and merge
Lastly, Figure 10 shows the wave topology “skeleton” of several example wave boundaries, where each branch of a wave is represented by a single line. These examples clearly highlight the complexity of some traffic waves, in which the wave merges and bifurcates several times as it propagates over several miles. Additionally, some wave boundaries have no branching at all, furhter suggesting the variability in complexity of different waves. A full analysis of typical wave topologies is outside of the scope of this paper; we only wish here to highlight the usefulness of a topology-based structure for clearly identifying complex wave phenomena such as merging and bifurcation.
@@
Figure 10: Demonstration of wave boundary topologies: thick blue lines represent the “trough” of the wave (where vehicles in the wave reach minimum speed). Thin blue “tie lines” indicate topology connections, i.e. a wave branching or merging. The gray outline is the overall wave boundary identified using the method from Section 4.3.4.

### v2

@@ 4 Methodology
NN the number of vehicle trajectories - -
@@
$\bm{\mathrm{\tau}}_{i}$ vector representing the trajectory of vehicle $i$ - -
$\bm{\mathrm{\tau}}_{i}^{c}$ vector representing the stationary component trajectory of vehicle $i$ - -
$\bm{\mathrm{\tau}}_{i}^{osc}$ vector representing the oscillation component trajectory of vehicle $i$ - -
@@
$d_{i}$ wave fronts set for vehicle $i$ - -
$a_{i}$ wave tails set for vehicle $i$ - -
$K_{i}$ the number of wave fronts and tails for vehicle $i$ - -
@@
$N_{c}$ the number of connected components for graph $\mathcal{G}$ - -
$\mathcal{C}^{d}$ the components for the wave fronts graph $\mathcal{G}_{d}$ - -
$\mathcal{C}^{a}$ the components for the wave tails graph $\mathcal{G}_{a}$ - -
@@ 4.1 Preliminaries
Long-range raw vehicle trajectory data as from [11] is imperfect due to the challenges of tracking and re-identification in dense traffic conditions [103; 104]. In essence, it is difficult to track vehicles perfectly across hundreds of cameras and through occlusions [105], so trajectory datasets over large spatial ranges often contain fragmented trajectories. These issues do not significantly impact local traffic measurements, such as traffic speeds, which can still be obtained with extremely high fidelity. We prepare raw data for the methods described next by utilizing the method from [106], which we describe in more detail in A. The resulting virtual trajectories have no passing or lane change maneuvers. This allows efficient search between adjacent trajectories, but the methods described next could also be applied to raw trajectory data with lane-change maneuvers with slight modifications.
@@ 4.2 Identifying the stop-and-go waves: the trajectory perspective
From the perspective of a vehicle trajectory, stop-and-go waves can be intuitively understood as a sequence of deceleration and acceleration events [16; 20]. This pattern is discernible from the speed time-series profile of each individual vehicle (see Figure 3). By analyzing these profiles, we can observe the fluctuations in speed that characterize stop-and-go waves, providing a clear visualization of how vehicles repeatedly slow down and speed up within the traffic flow.
@@
The identification process involves decomposing an empirical trajectory into two components [34], followed by detecting local maxima that represent wave fronts and minima that represent wave tails within the oscillation component. A flowchart with visual illustration of this process is provided in Fig. 4.
Figure 4: Flowchart (left) and visual demonstration (right) of wave identification for each vehicle trajectory: (a) speed profile with critical-speed threshold; (b) empirical vs. stationary component trajectory with vertical offset defining the oscillation component; (c) oscillation component time series with detected extrema.
Let $\mathcal{T}$ be the set of all trajectories. Note that the preprocessing in A ensures that trajectories are ordered by increasing time, and no trajectories enter or exit the considered lane. Let $\bm{\mathrm{\tau}}_{i}\in\mathcal{T}$ represent the trajectory of vehicle $i$ as a vector of time-space points. Throughout this section, we will use $\tau_{i}(t)$ to represent a continuous, functional representation of the trajectory of vehicle $i$ from which the points in $\bm{\mathrm{\tau}}_{i}$ are sampled. To detect the critical events from the trajectories, each trajectory is decomposed into two components as follows:
@@
By definition, a wave front or tail occurs when trajectory velocity $\dot{\tau}_{i}(t)=v_{c}$, or when $\dot{\tau}_{i}^{\text{osc}}(t)=0$. Thus, wave front and tail points correspond to a local maximum or minimum of $\tau_{i}^{osc}(t)$, respectively.
  • Wave front points. The wave front is associated with a local maximum of $\tau_{i}^{\text{osc}}(t)$, denoted by $t^{\text{front}}=t_{\text{local}}^{\text{max}}$.
  • Wave tail points. The wave tail is associated with a local minimum of $\tau_{i}^{\text{osc}}(t)$, denoted by $t^{\text{tail}}=t_{\text{local}}^{\text{min}}$.
In our implementation, the local maxima and minima of $\tau_{i}^{\text{osc}}(t)$ are identified using peak-finding algorithms scipy.signal.find_peaks [107]. We find all local maxima and minima in $\bm{\mathrm{\tau}}_{i}^{osc}$, yielding a set $d_{i}$ of wave front and a set $a_{i}$ of wave tail points for trajectory $i$. We keep only those wave fronts that are followed by a wave tail and those wave tails that are preceded by a wave front, excluding any wave fronts or tails that occur at the boundaries of the trajectories, as they do not contribute to a complete stop-and-go cycle. This ensures that there are the same number $K_{i}$ of wave front points and wave tail points for trajectory $i$. This process is repeated for all trajectories in $\mathcal{T}$, yielding an overall set of wave front and tail points $\mathcal{D}=\{d_{0},d_{1},...d_{i}\}$ and $\mathcal{A}=\{a_{0},a_{1},...a_{i}\}$.
@@ 4.3.2 Edges
Next, we describe the process for adding edges to the stop-and-go wave graph such that when finished, all connected wave fronts and wave tails correspond to a single distinct traffic wave. Two types of edges are considered to connect nodes in the stop-and-go-wave graph. i). Inner edges connect wave fronts and tails for the same trajectory, based on the intuition that each wave front and tail are added to the graph as a pair and by definition correspond to the same stop and go wave. ii.) Cross edges connect either wave fronts or wave tails for nearby trajectories, based on the idea two decelerations or accelerations experienced by two trajectories in close spatio-temporal proximity are caused by the same stop-and-go wave. This framework facilitates the analysis of both individual trajectory patterns and the interactions between different trajectories [16; 15].
@@
    $\mathcal{E}_{\text{inner}}=\bigcup_{i=1}^{N}\bigcup_{k_{i}=1}^{K_{i}}\{\{d_{i,k_{i}},a_{i,k_{i}}\}\}.$ (5)
@@
    $\mathcal{E}_{\text{cross}}=\mathcal{E}_{\text{cross}}^{\mathcal{D}}\cup\mathcal{E}_{\text{cross}}^{\mathcal{A}},$ (6)
    where the connection for wave front is illustrated by the red dashed lines connecting adjacent red dots in 5b, green for wave tails. Succinctly, for each wave front point, we search in a narrow spatio-temporal box in time and in space around that point for other wave front points identified on the next upstream trajectory. If any are found, we add an edge to $\mathcal{E}_{\text{cross}}^{\mathcal{D}}$. If there are multiple points found, it will connect to the one closest in time. The same is then done for each wave tail point. Mathematically, we define these sets of edges connecting wave fronts ($\mathcal{E}_{\text{cross}}^{\mathcal{D}}$) and tails ($\mathcal{E}_{\text{cross}}^{\mathcal{A}}$) as:
@@ 5 Data
I-24 MOTION is a traffic instrument for the freeway traffic observation [111; 11; 104], which is designed for the continuous freeway traffic data collection and analysis. It is expected to generate 200 million vehicle miles traveled (VMT) in trajectory data annually [104], which poses significant challenges for data analysis. In this paper, the I-24 MOTION INCEPTION data released by [11] is used for the stop-and-go wave analysis with the proposed method in Section 4. Table 4 summarizes the data used in this paper, including the corresponding week day of the date, vehicle miles traveled (VMT) and the mean speed (MS). The method input consists of virtual trajectories generated from the mean speed field, designed to address data quality issues [103] in individual trajectories and enable scalable measurements since traffic waves are the phenomena in between microscopic and macroscopic. Further details are provided in A.
@@
The hyper-parameter in this paper is the critical speed $v_{c}$ and is examined at various values: $\{1,5,10,15,20,25,30,35,40\}$ mph. The search region $\mathcal{S}$ as illustrated in Figure 5, is defined as a rectangular area centered around a given spatio-temporal feature, extending from -0.02 miles to 0.05 miles spatially and from -5 seconds to 15 seconds temporally. In this context, the negative direction in space refers to the opposite direction of traffic flow, while negative values in time indicate past moments relative to the spatio-temporal feature. The 15-second section is chosen based on the 5-second interval for sending virtual vehicles to the field, as outlined in A. The selected size should neither be excessively large nor too small; the criterion is to ensure adequate coverage of the next vehicle’s adjacent critical spatio-temporal point (either the front or tail). The design of looking backward in time and space is to account for specific cases where waves are stationary [112] or when drivers engage in preventive driving, resulting in braking points occurring even earlier than those of the preceding vehicle.
6 Results: Wave boundaries identification
@@ 6.2 Wave fronts and tails identification
For each $v_{c}$ and each lane, a set of wave fronts and tails is generated. Figures 8(a) and 8(b) illustrate the average number of wave fronts and tails identified per day, respectively. As illustrated in the figures, the number of wave fronts and tails increases as the critical speed $v_{c}$ rises. The number of wave fronts and tails tends to be higher in lane 4, followed by lane 3, lane 2, and lane 1, for speeds ranging from 10 to 40 mph. However, as shown in Figures 8(c) and 8(d), the traveling distance of the fronts and tails follows the inverse order, indicating that wave propagation is more prolonged in the inner lanes compared to the outer lanes, which may be explained by the boundary effects of on-ramps and off-ramps.
@@
Figures 8(c) and 8(d) demonstrate that the travel distances of wave fronts and tails exhibit nonlinear variations with the critical speed $v_{c}$ across four lanes. The travel distance reaches its maximum around a critical speed of 15-20 mph for all lanes, indicating that this speed range is most critical for wave propagation. Lane 1 consistently has the longest wave front travel distance, while lane 4 shows the shortest across all speed levels. These findings suggest that there is a critical speed range for maximizing wave propagation distance, with notable differences in wave dynamics between lanes, particularly at lower and higher speeds. This emphasizes the sensitivity of wave behavior to both lane-specific factors and the chosen critical speed.
@@ 6.3 Wave components identification
7 Results: Statistics of traffic waves dynamics
@@ 7.2 Differences in traveling speeds between wave fronts and tails
Based on the definitions in Equations (12) and (13), each sample (i.e., wave front or tail) can be represented in a scatter plot illustrating the relationship between wave travel speed and travel distance, as shown in Figure 10. Note that the speed shown here are absolute values; in reality, these speeds represent the traveling waves in the opposite direction of traffic flow. Figure 10 illustrates the centrality of wave speed for both the fronts and tails as the wave travel distance increases at the critical speed 15 mph. The wave travel speed for distances below 0.5 miles is highly random across all lanes. Additional plots for different critical speeds can be found in E.
@@ 7.3 Wave bifurcation and merge
Lastly, Figure 11 shows the wave topology “skeleton” of several example wave boundaries, where each branch of a wave is represented by a single line. These examples clearly highlight the complexity of some traffic waves, in which the wave merges and bifurcates several times as it propagates over several miles. Additionally, some wave boundaries have no branching at all, further suggesting the variability in complexity of different waves. We further discuss the clusters of the identified wave topologies in the next section.
@@
Figure 11: Demonstration of wave boundary topologies: thick blue lines represent the “trough” of the wave (where the middle point of the paired wave front and tail). Thin blue “tie lines” indicate topology connections, i.e. a wave branching or merging. The gray outline is the overall wave boundary identified using the method from Section 4.3.4.
7.4 Clustering analysis on the wave topology
Clustering analysis can help better understand the topological structure of the waves. We use a simple clustering method, KMeans clustering via scikit-learn [113], and visualize the results on the diagram shown in Figure 12. Figure 12 is a demonstration of lane 1 on the critical speed 15 mph. We cluster the wave components lane by lane since we think the topological features on each lane are different. The cluster visualization clearly shows the simplicity and complexity of traffic waves: traffic waves with similar travel distances may remain stable without any bifurcation or merging, while others may have bifurcation and merging more than 10 times.
(a) Bifurcation
(b) Merge
Figure 12: Cluster visualization for lane 1 with the critical speed set at 15 mph: The x-axis represents the number of bifurcations/merges of the wave components, and the y-axis represents the wave length of the wave component.
  • Cluster 1 in Figure 12 are waves with simple topological structures: the wave components in this cluster can travel between 0 and 4 miles, experiencing either no bifurcations and merges or only 1-2 occurrences. The wave topology remains relatively stable. Figure 13(c) demonstrate one sample topology from this cluster.
  • Cluster 2 falls in between, being neither highly complex nor overly simple. Figure 13(b) demonstrate one sample topology from this cluster.
  • Cluster 3 consists of complex topologies, where a single wave can bifurcate and merge back more than 10 times over a travel distance of more than 3 miles. Figure 13(a) demonstrate one sample topology from this cluster.
(a) (a) Sample topology from cluster 3
(b) (b) Sample topology from cluster 2
(c) (c) Sample topology from cluster 1
Figure 13: Topology samples from different clusters: (a) A sample wave topology from the most complex cluster bifurcates 17 times and merges 15 times, traveling 3.62 miles; (b) A sample wave topology from the intermediate complexity cluster bifurcates 5 times and merges 2 times, traveling 3.58 miles; (c) A sample wave topology from the simplest cluster has no bifurcations or merges, traveling 3.64 miles.
7.5 Spatial distribution of wave bifurcation and merge
With all the identified wave topologies, we further plot the spatial locations of wave bifurcation and merging points to examine their relationship with ramps. Figure 14 demonstrates the relationship between wave bifurcation (one wave split to multiple waves) and merge points (multiple waves combine to one) across lanes regarding their proximity (absolute distance) to the closest on/off-ramp points [11].
(a) Lane 1 wave bifurcation
(b) Lane 1 wave merge
(c) Lane 2 wave bifurcation
(d) Lane 2 wave merge
(e) Lane 3 wave bifurcation
(f) Lane 3 wave merge
(g) Lane 4 wave bifurcation
(h) Lane 4 wave merge
Figure 14: Wave bifurcation and merge points spatial distribution across all lanes
The histograms show the frequency distribution of wave bifurcations and merges occurrence at different distances from the on/off-ramps. The figures in the left column (shaded light blue) show the distribution of wave bifurcations, while the figures in the right column (shaded orange) show the distribution of wave merges.
From Figure 14, we can observe that wave bifurcation occurs closer to ramps, while wave merge happens slightly further from ramps, with a more balanced distribution. However, it does reveal a pattern indicating that road geometry features can explain the occurrence of wave bifurcation and merging, as they are more concentrated in the proximity distribution. On the other hand, wave bifurcation and merging also occur in areas located 0.5 miles away from any on-ramps or off-ramps, suggesting that driver behavior may also play a role in these complex dynamics.
Although identifying the root cause of bifurcation and merging is challenging, our analysis of road geometry suggests that these phenomena can occur anywhere, although with a concentration within the ramp impact area. We believe the underlying cause is more closely tied to driving behavior, with road geometry acting as a contributing factor by influencing the frequency of cut-ins and lane changes. Factors such as asymmetric driving behavior [114; 16], anticipation effects [115], and driver distraction [73] may play a significant role. A more detailed microscopic analysis of driving behavior, enabled by improvements in data quality [103], could provide deeper insights in the future.